OpenStudy (anonymous):

Slove cscx+2=0 0 <= x < 2pi

OpenStudy (anonymous):

2 csc+2 x -1 0 <= 0 2pt and 2 cos x-1=0 0 <= < 2pi please help!!!

OpenStudy (loser66):

csc +2 =0 ---> csc =-2 and csc = 1/ sin so, 1/sin = -2